OpenStudy (anonymous):

write the point- slope form of the equtaion of the line through the given point with the given slope . through : ( -5,-5) , slope = 1/5

OpenStudy (ajprincess):

if the given point is (x_1,y_1) and the slope is m then the point- slope form of the equtaion is given by y-y_1=m(x-x_1) Does that help?

OpenStudy (anonymous):

yea thanks

OpenStudy (ajprincess):

Welcome:)

OpenStudy (anonymous):

mind helping me with something else ?

OpenStudy (ajprincess):

ask me:)

OpenStudy (anonymous):

how do you write the slop - intercept form for an equation ?

OpenStudy (anonymous):

slope

OpenStudy (ajprincess):

y-y_1=m(x-x_1) y=m(x-x_1)+y_1->slope-intercept form

OpenStudy (ajprincess):

Does that help?